OpenStudy (anonymous):

Please help Which is the exact area of the circle? A circle is shown with its radius marked 9 feet eighty-one square feet eighty-one times pi square feet eighty-one times pi squared square feet eighty-one squared times pi squared square feet

OpenStudy (anonymous):



OpenStudy (xmoses1):

Use \[\pi~r^2\]

OpenStudy (anonymous):

i don't get that

OpenStudy (anonymous):

like war would i do

OpenStudy (anonymous):

what not war sorry

OpenStudy (xmoses1):

\( \pi \) * 9 squared. Tell me what \( {9}^2 \) is

OpenStudy (anonymous):

81

OpenStudy (xmoses1):

Ok and then you would multiply that times \( \pi \)

OpenStudy (anonymous):

so 81*PI square feet

OpenStudy (xmoses1):

so \( 81* \pi \) Yes that is correct

OpenStudy (anonymous):

ok thankyou

OpenStudy (xmoses1):

You are welcome :)